OpenStudy (anonymous):

1+3(5-17) diviede by 2.6 a. -4 c. -107 b. 109 d. -144

OpenStudy (anonymous):

Using PEMDAS, we have to distribute the 3(5-17), which, broken down, is 3 * 5 and 3 * -17.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1+15-51DIVIDE BYE 2.6

OpenStudy (anonymous):

So, your solution becomes 1 + (15) + (-51). So, 1+ 15 = 16. Plug this in, so our solution becomes 16 + (-51).

OpenStudy (anonymous):

WERE DID YOU GET THE 16

OpenStudy (anonymous):

The 1 + 15

OpenStudy (anonymous):

We had 1 in the original, and we had to times 3 by 5, which gives us 15. So, 1 + 15 = 16. Understand?

OpenStudy (anonymous):

OOK

OpenStudy (anonymous):

Is the original question:\[\frac{ 1 + 3(5-17) }{ 2.6 }\]

OpenStudy (anonymous):

YEA

OpenStudy (anonymous):

One moment.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

start by simplifying the number in the parenthesis

OpenStudy (anonymous):

5 - 17 = what?

OpenStudy (anonymous):

@billj5 Don't you have to distribute the 3 to the 5 and -17?

OpenStudy (anonymous):

no, either way it gives you same number

OpenStudy (anonymous):

Yeah, it does. I'm stumped, the answer I'm getting isn't any of the answers.

OpenStudy (anonymous):

you can do it that way though, it doesnt matter

OpenStudy (anonymous):

I'm getting\[\frac{ -35 }{ 2.6 } = -13.46\]

OpenStudy (anonymous):

im get in decimals

OpenStudy (anonymous):

Am I just doing something wrong?

OpenStudy (anonymous):

no something is wrong with the original question or the answers

OpenStudy (anonymous):

Has to be.

OpenStudy (anonymous):

the answers is -107 cuz i answer randomely and i got it right but i still want to know why is -107

OpenStudy (anonymous):

look at the question again, make sure its what Nibby asked it was

OpenStudy (anonymous):

the question is 1+3(5-17)/2.6

OpenStudy (lunymoony):

I believe the list of answers are all wrong, should be something like 1.7, but yeah, have to say C.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

have to say c because rocky1718 said it was correct or for some other reason?

OpenStudy (lunymoony):

If you grow up I will give you a medal

OpenStudy (anonymous):

@LunyMoony que?

OpenStudy (lunymoony):

que? @billj5 ..what u mean?

OpenStudy (anonymous):

@LunyMoony "if you grow up I will give you a medal? what you mean?

OpenStudy (lunymoony):

Sorry, didnt know what you meant but I believe no one likes it when people act immature and give out random drama. When I said it has to be C is because the problem I was doing rounded out to something like 1.7.., so we kind of have to assume its 107 they want. Again, sorry, for the misunderstanding @billj5 You are a good mathematician

OpenStudy (anonymous):

@LunyMoony its ok man, calm down, i really wanted to know what you meant, thank you for answering, next time maybe you wont get offended by someone asking a real question

OpenStudy (lunymoony):

Explain your question because I still dont think I apprehend what you are asking @billj5

OpenStudy (anonymous):

you answered my question when you said "When I said it has to be C is because the problem I was doing rounded out to something like 1.7.., so we kind of have to assume its 107 they want. " ty

OpenStudy (anonymous):

have a good one

OpenStudy (lunymoony):

you to @billj5